Find the curvature, the radius and the center of curvature at a point.


y=sin x, x = pi/2


1
Expert's answer
2021-12-06T16:39:24-0500
"K=\dfrac{|y''|}{[1+(y')^2]^{3/2}}"

"y'=\cos x, y''=-\sin x"

"K=\dfrac{|-\sin x|}{[1+\cos ^2 x]^{3/2}}"

"x=\pi/2"

"K=\dfrac{|-\sin (\pi/2)|}{[1+\cos ^2 (\pi/2)]^{3/2}}=1"



The radius of curvature of a curve at a point is called the inverse of the curvature "K" of the curve at this point:


"R=\dfrac{1}{K}=1"

"x_C=x-\dfrac{y'(1+(y')^2)}{y''}"

"y_C=y+\dfrac{1+(y')^2}{y''}"

"x=\pi/2"

"y(\pi/2)=\sin(\pi/2)=1"

"y'(\pi/2)=-\cos(\pi/2)=0"

"y''(\pi/2)=-\sin(\pi/2)=-1"

"x_C=\pi/2-\dfrac{0(1+(0)^2)}{-1}=\pi/2"

"y_C=1+\dfrac{1+(0)^2}{-1}=0"

"C(\pi/2, 0)"